Section 2.5-Constant-Coordinate Surfaces 2.31 Describe the intersection of the following surfaces: (a) * = 2, y = 5 (b) x = 2 2.31 solution
Answer #1

cal intersection is line Cary, z) = (215,2)= (215,0) + (0,0,1) 7 ro intersection is point (x, y, z) = (2,-1, 10) w intersecti© Zz o cosø = 10 . -id T = 10 rs 20; x² + y²+ z = 20² x?+ y + 10% = 202 re2+ y = 300 This is circle in plane, t=10 © = rcosø